# Strong Quantum Mpemba Effect with Squeezed Thermal Reservoirs

### v1

@@ Abstract
The phenomena where a quantum system can be exponentially accelerated to its stationary state has been refereed to as Quantum Mpemba Effect (QMpE). Due to its analogy with the classical Mpemba effect, hot water freezes faster than cold water, this phenomena has garnered significant attention. Although QMpE has been characterized and experimentally verified in different scenarios, sufficient and necessary conditions to achieve such a phenomenon are still under investigation. In this paper we address a sufficient condition for QMpE through a general approach for open quantum systems dynamics. With help of the Mpemba parameter introduced in this work to quantify how strong the QMpE can be, we discuss how our conditions can predict and explain the emergence of weak and strong QMpE in a robust way. As application, by harnessing intrinsic non-classical nature of squeezed thermal environments, we show how strong QMpE can be effectively induced when our conditions are met. Due to the thermal nature of environment considered in our model, our work demonstrates that a hot qubit freezes faster than a cold qubit only in presence of squeezed reservoirs. Our results provide tools and new insights opening a broad avenue for further investigation at most fundamental levels of this peculiar phenomena in the quantum realm.
1 Introduction
@@
Two physical systems initially at different temperatures, the cold $T_{\mathrm{c}}$ and hot $T_{\mathrm{h}}>T_{\mathrm{c}}$, are cooled down by exchanging energy with a colder reservoir at temperature $T_{0}<T_{\mathrm{c}}$. Even under the same cooling conditions, the hot system can reach regime of temperatures smaller than the cold system does during the process. This unconventional effect was observed in the classical physics realm by a 13 years old high school student, called Erasto B. Mpemba, in 1963. The experimental verification of the effect was reported by E. B. Mpemba and D. G. Osborne in 1969 [1]. F. Carollo et al [2] introduced a mechanism able to exponentially speed up the dynamics of a quantum system to the stationary state, which can be interpreted as a Mpemba’s effect in the quantum realm. Since then, such a quantum Mpemba’s effect (QMpE) has aroused further investigations about the phenomena and its characterization in different contexts [3, 4, 5, 6, 7, 8, 4, 9, 10, 11, 12, 13, 14].
Figure 1: (a) Two identical two-level systems are prepared initially in thermal states at different temperatures $T_{c}$ and $T_{h}$. The systems are cooled down independently to a target colder temperature $T_{0}<T_{c}$. During the cooling process the population of the excited states $|e\rangle$ of each system is transferred to the ground state $|g\rangle$. The emergence of QMpE depends on the squeezing parameter $r$ of the bath, where (b) QMpE is not observed in our model for Gaussian thermal baths ($r=0$), but (c) by considering non-zero squeezing parameters, the QMpE is engineered.
@@
The QMpE has been investigated in the context of Markovian dynamics of many-body systems through Dicke model [2], Anderson model (quantum dots reservoirs) [3], Lieb-Liniger model [4], and spin chain systems [5, 6] for example. However, to the best of our knowledge, the first experimental evidence of such an effect was done recently by J. Zhang et al [15] in a single trapped ion systems, exploiting the dynamics of tree-level system encoded in the low-lying electronic states of a trapped 40Ca+ ion. However, all these works, among others [8, 4, 9], have considered the interaction of the physical system with reservoirs described by physical processes other than that one observed in thermal reservoirs with well defined finite temperature ($T>0$). To deal with the proper measure and characterization, some authors [3, 10] have considered the definition of an alternative version of temperature for isothermal processes, where detailed study on the exchanged heat and entropy variation of the system is considered.
@@
As depicted in Fig. 1a, a genuine QMpE should be possible if the initial hot and cold states are prepared using the same reservoir at different temperatures, and then use the same bath at a colder temperature to induce the QMpE. In addition, despite efforts to observe and characterize the QMpE, the specific conditions required to induce this effect and the mechanism behind it remain to be explored in greater detail. In this regard, we address these two questions. Firstly, we conveniently exploit the theory of open quantum system in the superoperator formalism to derive fundamental conditions that should be satisfied to efficiently achieve QMpE [16]. Such conditions are then used to show that, in the context of reservoir, Hamiltonian and state preparation engineering, they provide a consistent way of explaining the mechanism behind the emergence of QMpE. With help of a fidelity-base figure of merit introduced in this work, as shown in Figs. 1b and 1c, we identify the emergence of thermally induced QMpE in a two-level atom interacting with squeezed thermal reservoirs. In addition to the application of squeezed in gravitational-wave detection [17, 18, 19], quantum metrology [20, 21], efficient quantum thermal engines [22, 23, 24, 25], among others [26], our work establishes the impact of squeezed thermal baths to reached out genuine QMpE.
2 Sufficient conditions for QMpE
@@
First of all, let us introduce a function (or quantifier) used as a figure of merit (or witness) of the Mpemba’s effect. Such a function is defined from the measure of the distance (or infidelity) between two quantum states, introduced in context of quantum information [27], obtained from the trace distance $\mathcal{B}_{\mathrm{c},\mathrm{h}}(t)=||\hat{\rho}_{\mathrm{c},\mathrm{h}}(t)-\hat{\rho}_{0}||_{1}$, with $||\hat{A}||_{1}=[\mathrm{Tr}(\hat{A}^{\dagger}\hat{A})]^{1/2}$. $\hat{\rho}_{\mathrm{c}}(t)$ and $\hat{\rho}_{\mathrm{h}}(t)$ are solutions of the Lindblad equation for the initial condition $\hat{\rho}_{\mathrm{c}}(0)=\hat{\rho}_{\mathrm{c}}$ and $\hat{\rho}_{\mathrm{h}}(0)=\hat{\rho}_{\mathrm{h}}$, respectively. From this equation, and with help of Figs. 1b and 1c, it is possible to state that whenever the initial states satisfy $\mathcal{B}_{\mathrm{h}}(0)>\mathcal{B}_{\mathrm{c}}(0)$, the Mpemba’s effect is observed when $\mathcal{B}_{\mathrm{h}}(t)<\mathcal{B}_{\mathrm{c}}(t)$, as $\mathcal{B}$ quantifies the distance to the steady state. Therefore, we define the infidelity-based quantity
@@
The function $\mathcal{A}_{\mathcal{B}_{\mathrm{c}}>\mathcal{B}_{\mathrm{h}}}$ quantifies the (time-averaged) contributions of the fidelity when the hot state gets closer to the target state in comparison to the colder one (as sketched in Fig. 1b). On the other hand, $\mathcal{A}_{\mathcal{B},0}$ quantifies the time average of the absolute value of the difference between the $\mathcal{B}_{\mathrm{c}}(t)$ and $\mathcal{B}_{\mathrm{h}}(t)$ during the total evolution time $\tau$. By construction, $\mathcal{M}$ has to be smaller than $1$, as the initial condition of our problem we have to choose quantum states such that $\mathcal{B}_{\mathrm{c}}(0)<\mathcal{B}_{\mathrm{h}}(0)$, as depicted in Fig. 1. Also, it possible to observe that the integral function will capture any contribution of the dynamics in which $\mathcal{A}_{\mathcal{B}_{\mathrm{c}}>\mathcal{B}_{\mathrm{h}}}$, such that $\mathcal{M}>0$ indicates the emergence of the Mpemba’s effect within the integration interval considered. Since the absence of Mpemba’s effect leads to $\mathcal{M}=0$, we can conclude that our parameter obeys $10\leq\mathcal{M}<$ 1. In this way, the bigger the parameter $\mathcal{M}$, the stronger the Mpemba’s effect.
@@
where $\hat{\rho}(t)$ is the system master equation and $\mathfrak{L}[\bullet]$ is the dynamics generator in the system’s Hilbert space $\mathcal{H}$, which encodes the information about the dynamics when the system interacts with external fields and its surroundings, for example. One alternative way to rewrite the above equation, and convenient to our discussion, is through the superoperator formalism [16, 28, 29, 30, 31] (see A for further details). In fact, through this formalism the Eq. (2) can be rewritten in a “Schrödinger form” as [16]
@@
where $|\varrho(t)\rangle\rangle$ is the coherence vector associated to the density matrix $\hat{\rho}(t)$ with components given by $\varrho_{n}(t)=\mathrm{tr}[\rho(t)\sigma^{\dagger}_{n}]$, for a basis $\{\sigma^{\dagger}_{n}\}$ of $D^{2}$ operators $\sigma^{\dagger}_{n}\in\mathcal{H}$, with $D=\dim[\mathcal{H}]$. Similarly, using the same basis we can define $\mathbbm{L}$ as a $D^{2}\times D^{2}$ operator with matrix elements $\mathbbm{L}_{kn}=\mathrm{tr}[\sigma_{k}^{\dagger}\mathfrak{L}[\sigma_{n}]]$, the super-operator Lindbladian. In this way, it is possible to observe that any information about the dynamics generator and the system state are now encoded in $\mathbbm{L}$ and $|\varrho(t)\rangle\rangle$, respectively. As consequence of this formalism, we can write the solution for the system dynamics as $|\varrho(t)\rangle\rangle=e^{\mathbbm{L}t}|\varrho(0)\rangle\rangle$. This formalism is convenient to our discussion because we can write the solution $|\varrho(t)\rangle\rangle$ as a linear decomposition of the left- and right-eigenvectors of $\mathbbm{L}$. In fact, without loss of generality, let us assume the set of left- and right-eigenvectors of $\mathbbm{L}$ satisfying $\mathbbm{L}|\mathcal{R}_{n}\rangle\rangle=\lambda_{n}|\mathcal{R}_{n}\rangle\rangle$ and $\langle\langle\mathcal{L}_{n}|\mathbbm{L}=\lambda_{n}\langle\langle\mathcal{L}_{n}|$, respectively, then one gets (see B for further details)
@@
with coefficients $\gamma_{n}=\langle\langle\mathcal{L}_{n}|\varrho(0)\rangle\rangle$. Therefore, we have now shown that the dynamics of the coherence vector is uniquely obtained by the initial state and the spectrum of $\mathbbm{L}$. In this way, we can now study the expected behavior of the system in the steady-state, as well as the relevant decay rates of interest to emergence of the Mpemba effect.
Figure 2: Mpemba parameter $\mathcal{M}_{\mathcal{B}}$ as function of the squeezing degree $r$ and the external driving phase $\varphi$ for (a) $\Omega/\Delta_{0}=0.5$ and (b) $\Omega/\Delta_{0}=1.0$. The distance between the hot and cold states with respect to the coldest steady state is shown in (c,d,e) for the highlighted points I, II and III of (a), respectively. Similarly, the distance shown in (c,d,e) is related to the highlighted points I, II and III of (b), respectively. The temperature for the initial states is considered as $\hbar\omega_{0}/T_{c}k_{B}=2.0$ and $\hbar\omega_{0}/T_{h}k_{B}=0.1$ for the cold and hot states, and $\hbar\omega_{0}/T_{0}k_{B}=100$ for the coldest target state. We set the squeezing phase $\varphi=0$.
First of all, it is worth to state that the classical and quantum versions of the Mpemba effect only depends on the initial states of the hot and cold system, and the common environment used to cooling down the systems independently. Within the framework considered here for the QMpE, this means that the hot and cold systems have initial states represented by different coherence vectors, say $|\varrho_{\mathrm{c}}(0)\rangle\rangle$ and $|\varrho_{\mathrm{h}}(0)\rangle\rangle$, respectively, but they are driven through the same super-operator Lindbladian, and therefore we can write down
@@
where we already separated the sum into two parts associated to all eigenvalues $\lambda_{n}$ with $\mathrm{Re}[\lambda_{n}]=0$, and $\mathrm{Re}[\lambda_{n}]\neq 0$. The physical intuition behind this distinction relies on the fact that the first term of the above equation may describe the existence of one or multiple steady states, and the second term vanishes for sufficiently large evolution time.
@@
In context of QMpE, the main conclusion of Eq. (5) can be interpreted as follows. By ordering the eigenvalues such that $\mathrm{Re}[\lambda_{0}]\geq\mathrm{Re}[\lambda_{1}]\geq\cdots\geq\mathrm{Re}[\lambda_{D^{2}-1}]$, the component of $|\varrho_{\mathrm{c}/\mathrm{h}}(t)\rangle\rangle$ associated to $\mathrm{Re}[\lambda_{D^{2}-1}]$ will decay much faster than the component of $\mathrm{Re}[\lambda_{1}]$. It means that a hot state obeying $|\gamma^{\mathrm{h}}_{1}|\leq|\gamma^{\mathrm{h}}_{2}|\leq\cdots\leq|\gamma^{\mathrm{h}}_{D^{2}-1}|$ should approaches the steady state faster than an cold state with $|\gamma^{\mathrm{c}}_{1}|\geq|\gamma^{\mathrm{c}}_{2}|\geq\cdots\geq|\gamma^{\mathrm{c}}_{\lambda_{D^{2}-1}}|$. In addition to these conditions, it is also relevant to observe the relation between the coefficients associated to the eigenvalue most negative real part $\lambda_{D^{2}-1}$ and the first eigenvalue with a non-zero real part, say $\lambda_{x}$, such that $\mathrm{Re}[\lambda_{x}]<0$. According to Eq. (5), one expects to be able to observe strong QMpE if the additional relations $|\gamma^{\mathrm{h}}_{\lambda_{D^{2}-1}}|\geq|\gamma^{\mathrm{c}}_{\lambda_{D^{2}-1}}|$ and $|\gamma^{\mathrm{h}}_{\lambda_{x}}|\leq|\gamma^{\mathrm{c}}_{\lambda_{x}}|$, between the cold and hot initial states is satisfied.
Notice that, differently from Ref. [4], we are not providing conditions to characterize QMpE (this is done through the parameter $\mathcal{M}$), the inequalities stated are expected to be satisfied for any hot and cold states in which the QMpE can be reached out. Of course, since an analytical solution of $|\varrho_{\mathrm{c}/\mathrm{h}}(t)\rangle\rangle$ for an arbitrary dynamics is impossible, the degree of rigor of our approach only allows to state that these conditions are necessary, but not sufficient for QMpE. It means that, the absence of QMpE can be observed even when satisfying these inequalities, but it is not expected to observe QMpE if such conditions are not satisfied. Now, the relevance and applicability of these conditions will be discussed in a concrete example.
3 QMpE with squeezed thermal baths

### v2

# Enhanced Quantum Mpemba Effect with Squeezed Thermal Reservoirs

@@ Abstract
The phenomenon where a quantum system can be exponentially accelerated to its stationary state has been referred to as the Quantum Mpemba Effect (QMpE). Due to its analogy with the classical Mpemba effect, hot water freezes faster than cold water, this phenomenon has garnered significant attention. Although QMpE has been characterized and experimentally verified in different scenarios, the sufficient and necessary conditions to achieve such a phenomenon are still under investigation. In this paper, we address a sufficient condition for QMpE through a general approach for open quantum system dynamics. With the help of the Mpemba parameter introduced in this work to quantify how strong the QMpE can be, we discuss how our conditions can predict and explain the emergence of weak and strong QMpE in a robust way. As an application, by harnessing the intrinsic non-classical nature of squeezed thermal environments, we show how enhanced QMpE can be effectively induced when our conditions are met. We demonstrate that when the system interacts with thermal reservoirs, a hot qubit freezes faster than a cold qubit in the presence of squeezing. Our results provide tools and new insights, opening a broad avenue for further investigation at the most fundamental levels of this peculiar phenomenon in the quantum realm.
I Introduction
@@
Two physical systems initially at different temperatures, the cold $T_{\mathrm{c}}$ and hot $T_{\mathrm{h}}>T_{\mathrm{c}}$, are cooled down by exchanging energy with a colder reservoir at temperature $T_{0}<T_{\mathrm{c}}$. Even under the same cooling conditions, the hot system can reach a regime of temperatures smaller than those the cold system does during the process. This unconventional effect was observed in the classical physics realm by a 13-year-old high school student, Erasto B. Mpemba, in 1963. The experimental verification of the effect was reported by E. B. Mpemba and D. G. Osborne in 1969 Mpemba:1969 . F. Carollo et alFedericoCarolo:21 introduced a mechanism capable of exponentially speeding up the dynamics of a quantum system to the stationary state, which can be interpreted as Mpemba’s effect in the quantum realm. Since then, this quantum Mpemba effect (QMpE) has aroused further investigations about the phenomenon and its characterization in different contexts Chatterjee:23 ; Rylands:23 ; Kochsiek:22 ; Ares:23 ; PhysRevLett.133.140404 ; Manikandan:21 ; Rylands:23 ; Ivander:23 ; Chatterjee:23-b ; Liu:24 ; Turkeshi:24 ; Murciano:24 ; Yamashika:24 .
@@
The QMpE has been investigated in the context of Markovian dynamics of many-body systems through the Dicke model FedericoCarolo:21 , the Anderson model (quantum dots reservoirs) Chatterjee:23 , the Lieb-Liniger model Rylands:23 , and spin chain systems Kochsiek:22 ; Ares:23 , for example. However, to the best of our knowledge, the first experimental evidence of such an effect was provided recently by J. Zhang et alZhang:24 in a single trapped ion system, exploiting the dynamics of a three-level system encoded in the low-lying electronic states of a trapped 40Ca+ ion. However, all these works, among others Manikandan:21 ; Rylands:23 ; Ivander:23 , have considered the interaction of the physical system with reservoirs described by physical processes other than those observed in thermal reservoirs with a well-defined finite temperature ($T>0$). To deal with the proper measure and characterization, some authors Chatterjee:23 ; Chatterjee:23-b have considered the definition of an alternative version of temperature for isothermal processes, where a detailed study on the exchanged heat and entropy variation of the system is considered.
@@
As depicted in Fig. 1a, a genuine QMpE should be possible if the initial hot and cold states are prepared using the same reservoir at different temperatures, and then the same bath at a colder temperature is used to induce the QMpE. In addition, despite efforts to observe and characterize the QMpE, the specific conditions required to induce this effect and the mechanism behind it remain to be explored in greater detail. In this regard, we address these two questions. Firstly, we conveniently exploit the theory of open quantum systems in the superoperator formalism to derive fundamental conditions that should be satisfied to efficiently achieve QMpE Sarandy:05-2 . Such conditions are then used to show that, in the context of reservoir, Hamiltonian, and state preparation engineering, they provide a consistent way of explaining the mechanism behind the emergence of QMpE. With the help of a fidelity-based figure of merit introduced in this work, as shown in Figs. 1b and 1c, we identify the emergence of thermally induced QMpE in a two-level atom interacting with squeezed thermal reservoirs. In addition to the application of squeezed states in gravitational-wave detection McKenzie:02 ; Vahlbruch:10 ; Ligo:11 , quantum metrology Anisimov:10 ; Steinlechner:13 , efficient quantum thermal engines Klaers:17 ; Lindenfels:19 ; Xiao:23 ; Niedenzu:18 , among others Schnabel:17 , our work establishes the impact of squeezed thermal baths in achieving genuine QMpE.
Figure 1: (a) Schematic representation of the initial setup: two identical two-level quantum systems are prepared in thermal states at different initial temperatures, $T_{c}$ (cold) and $T_{h}$ (hot). Each system is then independently cooled down to a common target temperature $T_{0}<T_{c}$, with population transfer from the excited state $|e\rangle$ to the ground state $|g\rangle$ during the cooling process. (b) Time evolution of the systems in contact with a standard (unsqueezed) Gaussian thermal bath ($r=0$), where the quantum Mpemba effect (QMpE) is absent—that is, the initially hotter system does not relax faster than the colder one. (c) In contrast, the introduction of a non-zero squeezing parameter $r>0$ in the bath allows the engineering of QMpE. A highlighted region marks the regime where the initially hotter system relaxes more quickly than the colder one, a signature of the Mpemba effect emerging due to quantum bath squeezing.
II Sufficient conditions for QMpE
@@
First of all, let us introduce a function (or quantifier) used as a figure of merit (or witness) of Mpemba’s effect. Such a function is defined from the measure of the distance (or infidelity) between two quantum states, introduced in the context of quantum information Nielsen:Book , obtained from the trace distance $\mathcal{B}_{\mathrm{c},\mathrm{h}}(t)=||\hat{\rho}_{\mathrm{c},\mathrm{h}}(t)-\hat{\rho}_{0}||_{1}$, with $||\hat{A}||_{1}=[\mathrm{Tr}(\hat{A}^{\dagger}\hat{A})]^{1/2}$. $\hat{\rho}_{\mathrm{c}}(t)$ and $\hat{\rho}_{\mathrm{h}}(t)$ are solutions of the Lindblad equation for the initial conditions $\hat{\rho}_{\mathrm{c}}(0)=\hat{\rho}_{\mathrm{c}}$ and $\hat{\rho}_{\mathrm{h}}(0)=\hat{\rho}_{\mathrm{h}}$, respectively. From this equation, and with the help of Figs. 1b and 1c, it is possible to state that whenever the initial states satisfy $\mathcal{B}_{\mathrm{h}}(0)>\mathcal{B}_{\mathrm{c}}(0)$, Mpemba’s effect is observed when $\mathcal{B}_{\mathrm{h}}(t)<\mathcal{B}_{\mathrm{c}}(t)$, as $\mathcal{B}$ quantifies the distance to the steady state. Therefore, we define the infidelity-based quantity
@@
The function $\mathcal{A}_{\mathcal{B}_{\mathrm{c}}>\mathcal{B}_{\mathrm{h}}}$ quantifies the (time-averaged) contributions of the fidelity when the hot state gets closer to the target state in comparison to the colder one (as sketched in Fig. 1b). On the other hand, $\mathcal{A}_{\mathcal{B},0}$ quantifies the time average of the absolute value of the difference between $\mathcal{B}_{\mathrm{c}}(t)$ and $\mathcal{B}_{\mathrm{h}}(t)$ during the total evolution time $\tau$. To clarify the role of the total evolution time $\tau$ in the definition of $\mathcal{M}_{\mathcal{B}}$, notice that the factor $1/\tau$ appears in both the numerator and the denominator to ensure that $\mathcal{A}_{\mathcal{B}_{\mathrm{c}}>\mathcal{B}_{\mathrm{h}}}$ and $\mathcal{A}_{\mathcal{B},0}$ are dimensionless quantities, facilitating a proper comparison of the dynamics under different conditions. This normalization does not affect the qualitative interpretation of $\mathcal{M}_{\mathcal{B}}$, and in the asymptotic limit $\tau\rightarrow\infty$—corresponding to the steady state limit—the definition becomes independent of $\tau$, as both integrals converge. Therefore, the choice of including $\tau$ in Eq. (1) is made without loss of generality and is primarily for dimensional consistency.
By construction, $\mathcal{M}$ has to be smaller than $1$, since the initial condition of our problem requires us to choose quantum states such that $\mathcal{B}_{\mathrm{c}}(0)<\mathcal{B}_{\mathrm{h}}(0)$, as depicted in Fig. 1. Also, it is possible to observe that the integral function will capture any contribution of the dynamics in which $\mathcal{A}_{\mathcal{B}_{\mathrm{c}}>\mathcal{B}_{\mathrm{h}}}$, such that $\mathcal{M}>0$ indicates the emergence of Mpemba’s effect within the integration interval considered. Since the absence of Mpemba’s effect leads to $\mathcal{M}=0$, we can conclude that our parameter obeys $10\leq\mathcal{M}<$ 1. In this way, the bigger the parameter $\mathcal{M}$, the stronger Mpemba’s effect.
It is worth emphasizing that the Mpemba parameter $\mathcal{M}_{\mathcal{B}}$, introduced in Eq. (1), does not correspond directly to a relaxation rate or a specific physical observable. Rather, it is constructed as a mathematical tool to quantify the presence and relative strength of the quantum Mpemba effect. Similar to indicators used in non-Markovian dynamics Breuer:09 ; Rivas:10 ; Passos:19 , $\mathcal{M}_{\mathcal{B}}$ captures a global property of the dynamics: in this case, the extent to which the relaxation of the hot state is anomalously faster than that of the cold state over a finite time interval. This provides a practical and scalable approach to detect QMpE in both analytical models and numerical simulations, even when its direct operational interpretation is not accessible.
@@
where $\hat{\rho}(t)$ is the density matrix describing the system’s evolution, and $\mathfrak{L}[\bullet]$ is the dynamics generator in the system’s Hilbert space $\mathcal{H}$, which encodes the information about the dynamics when the system interacts with external fields and its surroundings, for example. An alternative way to rewrite the above equation, which is convenient to our discussion, is through the superoperator formalism Sarandy:05-2 ; Sarandy:05-1 ; Santos:20d ; Alicki:Book07 ; Petruccione:Book (see A for further details). In fact, through this formalism, Eq. (2) can be rewritten in a “Schrödinger form” as Sarandy:05-2 .
@@
where $|\varrho(t)\rangle\rangle$ is the coherence vector associated with the density matrix $\hat{\rho}(t)$ with components given by $\varrho_{n}(t)=\mathrm{tr}[\rho(t)\sigma^{\dagger}_{n}]$, for a basis $\{\sigma^{\dagger}_{n}\}$ of $D^{2}$-dimensional operators $\sigma^{\dagger}_{n}\in\mathcal{H}$, with $D=\dim[\mathcal{H}]$. Similarly, using the same basis, we can define $\mathbbm{L}$ as a $D^{2}\times D^{2}$ operator with matrix elements $\mathbbm{L}_{kn}=\mathrm{tr}[\sigma_{k}^{\dagger}\mathfrak{L}[\sigma_{n}]]$, the super-operator Lindbladian. In this way, it is possible to observe that any information about the dynamics generator and the system state is now encoded in $\mathbbm{L}$ and $|\varrho(t)\rangle\rangle$, respectively. As a consequence of this formalism, we can write the solution for the system dynamics as $|\varrho(t)\rangle\rangle=e^{\mathbbm{L}t}|\varrho(0)\rangle\rangle$. This formalism is convenient for our discussion because we can write the solution $|\varrho(t)\rangle\rangle$ as a linear decomposition of the left- and right-eigenvectors of $\mathbbm{L}$. In fact, without loss of generality, let us assume the set of left- and right-eigenvectors of $\mathbbm{L}$ satisfying $\mathbbm{L}|\mathcal{R}_{n}\rangle\rangle=\lambda_{n}|\mathcal{R}_{n}\rangle\rangle$ and $\langle\langle\mathcal{L}_{n}|\mathbbm{L}=\lambda_{n}\langle\langle\mathcal{L}_{n}|$, respectively, then we obtain (see B for further details).
@@
with coefficients $\gamma_{n}=\langle\langle\mathcal{L}_{n}|\varrho(0)\rangle\rangle$. Therefore, we have now shown that the dynamics of the coherence vector is uniquely determined by the initial state and the spectrum of $\mathbbm{L}$. In this way, we can now study the expected behavior of the system in the steady-state, as well as the relevant decay rates of interest for the emergence of the Mpemba effect.
Figure 2: (a, b) Mpemba parameter $\mathcal{M}_{\mathcal{B}}$ as a function of the squeezing parameter $r$ and the external driving phase $\varphi$, for two different ratios of system parameters: $\Omega/\Delta_{0}=0.5$ in (a) and $\Omega/\Delta_{0}=1.0$ in (b). This parameter quantifies the occurrence of the quantum Mpemba effect (QMpE), with regions of nonzero $\mathcal{M}_{\mathcal{B}}$ indicating where the initially hotter system cools faster than the colder one. (c, d, e) Time evolution of the trace distances $\mathcal{B}_{\mathrm{c}}(t)$ and $\mathcal{B}_{\mathrm{h}}(t)$, which measure the distinguishability between the instantaneous states (cold and hot initial conditions) and the coldest steady state. These are shown for the specific parameter sets labeled I, II, and III in panel (a), respectively. Similarly, the same analysis is performed in (f, g, h) for the corresponding highlighted points I, II, and III in panel (b). The initial temperatures are set as $\hbar\omega_{0}/(k_{B}T_{c})=2.0$ and $\hbar\omega_{0}/(k_{B}T_{h})=0.1$ for the cold and hot states, respectively, while the coldest target steady state corresponds to $\hbar\omega_{0}/(k_{B}T_{0})=100$. The squeezing phase is fixed at $\varphi=0$.
First of all, it is worth mentioning that the classical and quantum versions of the Mpemba effect depend only on the initial states of the hot and cold systems, and the common environment used to cool down the systems independently. Within the framework considered here for the QMpE, this means that the hot and cold systems have initial states represented by different coherence vectors, say $|\varrho_{\mathrm{c}}(0)\rangle\rangle$ and $|\varrho_{\mathrm{h}}(0)\rangle\rangle$, respectively, but they are driven through the same super-operator Lindbladian, and therefore we can write down
@@
where we already separated the sum into two parts associated with all eigenvalues $\lambda_{n}$ with $\mathrm{Re}[\lambda_{n}]=0$, and $\mathrm{Re}[\lambda_{n}]\neq 0$. The physical intuition behind this distinction relies on the fact that the first term of the above equation may describe the existence of one or multiple steady states, and the second term vanishes for sufficiently large evolution times.
@@
In the context of QMpE, the main conclusion of Eq. (5) can be interpreted as follows. By ordering the eigenvalues such that $\mathrm{Re}[\lambda_{0}]\geq\mathrm{Re}[\lambda_{1}]\geq\cdots\geq\mathrm{Re}[\lambda_{D^{2}-1}]$, the component of $|\varrho_{\mathrm{c}/\mathrm{h}}(t)\rangle\rangle$ associated with $\mathrm{Re}[\lambda_{D^{2}-1}]$ will decay much faster than the component associated with $\mathrm{Re}[\lambda_{1}]$. This means that a hot state obeying $|\gamma^{\mathrm{h}}_{1}|\leq|\gamma^{\mathrm{h}}_{2}|\leq\cdots\leq|\gamma^{\mathrm{h}}_{D^{2}-1}|$ should approach the steady state faster than a cold state with $|\gamma^{\mathrm{c}}_{1}|\geq|\gamma^{\mathrm{c}}_{2}|\geq\cdots\geq|\gamma^{\mathrm{c}}_{\lambda_{D^{2}-1}}|$. In addition to these conditions, it is also relevant to observe the relation between the coefficients associated with the eigenvalue with the most negative real part $\lambda_{D^{2}-1}$ and the first eigenvalue with a non-zero real part, say $\lambda_{x}$, such that $\mathrm{Re}[\lambda_{x}]<0$. According to Eq. (5), one expects to be able to observe an enhanced QMpE if the additional relations $|\gamma^{\mathrm{h}}_{\lambda_{D^{2}-1}}|\geq|\gamma^{\mathrm{c}}_{\lambda_{D^{2}-1}}|$ and $|\gamma^{\mathrm{h}}_{\lambda_{x}}|\leq|\gamma^{\mathrm{c}}_{\lambda_{x}}|$, between the cold and hot initial states are satisfied.
Notice that, differently from Ref. Rylands:23 , we are not providing conditions to characterize QMpE (this is done through the parameter $\mathcal{M}$); rather, the inequalities stated are expected to be satisfied for any hot and cold states in which the QMpE can be achieved. Naturally, obtaining an analytical solution for $|\varrho_{\mathrm{c}/\mathrm{h}}(t)\rangle\rangle$ under arbitrary dynamics is, in general, not feasible. As such, the level of analytical control in our approach only permits us to claim that the derived conditions are necessary, but not sufficient, for the occurrence of the quantum Mpemba effect (QMpE). Specifically, these conditions must be satisfied for QMpE to emerge within our model, but their fulfillment alone does not ensure the effect will manifest. In contrast, a violation of these conditions is sufficient to rule out the possibility of observing QMpE. This distinction is important: even if the inequalities are satisfied, other factors—such as interference between decay modes or coherence induced by squeezing—may prevent the effect from occurring. Therefore, the proposed conditions should be interpreted as rigorous filters that exclude the QMpE when not met, but do not guarantee its presence when they are. In the following, we illustrate the relevance and applicability of these conditions through a concrete example.
III QMpE with squeezed thermal baths